Given : TC=Q2+2Q+2 and TR=8Q-3 then
A. calculate profit maximizing level of output
B. maximum profit
1
Expert's answer
2019-12-18T08:59:12-0500

Solution:

Profit is TR minus TC

FWe compose the profit function and find its derivative:


"Profit=TR-TC"

"Profit=8Q-3-Q^2-2Q-2"

"Profit=-Q^2+6Q-5"

"Profit'=-2Q+6"

"-2Q+6=0"

"Q=3"

Answer: profit maximizing level of output  Q=3

B.


"Profit=-3^2+6\times 3-5=4"

Answer: maximum profit 4